# The Right to be Forgotten in Pruning: Unveil Machine Unlearning on Sparse Models

### v1

@@ 6 Related Work
Pruning. As the size of deep learning models increases, model pruning has attracted attention to reduce the parameter and improve computational efficiency without compromising accuracy [61, 62, 63, 64, 65, 66]. Existing pruning technologies can be classified into two types: structured pruning and unstructured pruning. In general, structured pruning [67, 68, 69, 70, 71, 72, 73, 74, 75, 76] remove sparse patterns such as layers and channels from the full model. Unstructured pruning [77, 78, 79, 80, 81, 82, 83] removes irregularly distributed parameters instead of entire weight structures (e.g., channels). Recent study shows that there is a dependence between pruned and the training data. Specific masks are more sensitive to the training data [60]. [9] proves that model Sparsitization can simplify the machine unlearning. Experiments by [84] conclude that the influence of deleted data on the connections between layers of the model remains unchanged.
Machine unlearning. Machine unlearning aims to remove the impact of deleted on trained models. While the retraining strategy is intuitive and costly, exact unlearning aims to learn a model with the same performance as the retrained model [6, 53, 85, 86, 87, 88, 89, 90, 91, 92, 93]. Approximate unlearning methods update the model by approximating the impact the deleted data on the training model [94, 95, 46, 96, 97, 98, 90, 99, 100]. Despite the success of these unlearning algorithms on dense models, unlearning on sparse models has not been extensively investigated. To the best the author’s knowledge,  [45] has also studied unlearning on sparse models. However, this work does not update the mask during the unlearning process (i.e., the right to be forgotten in pruning has not been addressed).
@@ Discussion
Structured [54] ResNet-20 Cifar10 1e-3 42 20%
Unstructured [101] ResNet-18 Cifar10 1e-3 42 20%
Unstructured [101] ResNet-18 FashionMNIST 1e-3 42 20%

### v2

@@ 6 Related Work
Pruning. As the size of deep learning models increases, model pruning has attracted attention to reduce the parameter and improve computational efficiency without compromising accuracy [DBLP:conf/nips/CunDS89, DBLP:journals/ijon/LiangGWSZ21, yin2023outlier, kumar2024no, wangexploring, bai2024sparsellm]. Existing pruning technologies can be classified into two types: structured pruning and unstructured pruning. In general, structured pruning [DBLP:conf/iccv/HeZS17, DBLP:conf/iccv/LiuLSHYZ17, DBLP:conf/iclr/0022KDSG17, DBLP:journals/corr/HuPTT16, DBLP:conf/nips/WenWWCL16, DBLP:conf/hpdc/HongSBKKNSC0S18, DBLP:conf/iclr/NonnenmacherPSR22, DBLP:conf/nips/HalabiSL22, DBLP:conf/aaai/YinUSJ023, DBLP:conf/icml/NovaDS23] remove sparse patterns such as layers and channels from the full model. Unstructured pruning [han2015deep, DBLP:conf/iclr/LiuSZHD19, DBLP:conf/fpga/CaoZYXNZLWZ19, DBLP:conf/asplos/RenZYLXQLW19, DBLP:conf/eccv/ZhangYZTWFW18, DBLP:conf/iclr/AlizadehTZAFLG22, DBLP:journals/corr/abs-2308-06619] removes irregularly distributed parameters instead of entire weight structures (e.g., channels). Recent study shows that there is a dependence between pruned and the training data. Specific masks are more sensitive to the training data [zhou2019deconstructing]. [liu2024model] proves that model Sparsitization can simplify the machine unlearning. Experiments by [zhang2022all] conclude that the influence of deleted data on the connections between layers of the model remains unchanged.
Machine unlearning. Machine unlearning aims to remove the impact of deleted on trained models. While the retraining strategy is intuitive and costly, exact unlearning aims to learn a model with the same performance as the retrained model [bourtoule2021machine, huang2024unified, yao2023large, zhao2024makes, DBLP:conf/aistats/LiWC21, DBLP:journals/corr/abs-2106-15093, DBLP:conf/icml/BrophyL21, DBLP:conf/nips/GinartGVZ19, cao2015towards, DBLP:conf/colt/UllahM0RA21, DBLP:conf/ccs/Chen [^0] HZ21]. Approximate unlearning methods update the model by approximating the impact the deleted data on the training model [chiencertified, park2024direct, DBLP:conf/aaai/GravesNG21, DBLP:conf/aaai/MarchantRA22, DBLP:conf/icml/WuDD20, DBLP:conf/alt/Neel0S21, DBLP:conf/nips/GinartGVZ19, DBLP:conf/aaai/FosterSB24, DBLP:conf/aaai/ChundawatTMK23]. Despite the success of these unlearning algorithms on dense models, unlearning on sparse models has not been extensively investigated. To the best the author’s knowledge,  [DBLP:conf/nips/JiaLRYLLSL23] has also studied unlearning on sparse models. However, this work does not update the mask during the unlearning process (i.e., the right to be forgotten in pruning has not been addressed).
@@ Discussion
Structured [he2018soft] ResNet-20 Cifar10 1e-3 42 20%
Unstructured [frankle2018lottery] ResNet-18 Cifar10 1e-3 42 20%
Unstructured [frankle2018lottery] ResNet-18 FashionMNIST 1e-3 42 20%